OpenStudy (anonymous):

If A is an n * n matrix such that AA^T = I, prove that det(A) = + or -1

OpenStudy (anonymous):

take det on both side det (A) * det(A^T)=det(I) det(A)=det(A^T), and det(I)=1 so det(A)^2=1 therefore det(A)=+ or -1